# A Backbone for Long-Horizon Robot Task Understanding

### v2

@@ III-C ActionREG: SAM-driven action registration network
A cornerstone of our TBBF is the ActionREG, designed for reasoning about and regressing object configurations (Fig. 5). This innovative component integrates therbligs’ prior knowledge into the SAM model as prompts, enabling sophisticated reasoning about the objects involved in robotic tasks. The integration of this knowledge serves as a guide for the model, improving its ability to understand and predict the configurations of objects within a task-specific context.
Algorithm 2 ActionREG and Context Matching
@@
0:  Object positions and orientations $\mathbf{P}_{\text{new}},\mathbf{O}_{\text{new}}$
@@
  return  $\mathbf{P}_{\text{new}},\mathbf{O}_{\text{new}}$
Through ActionREG, we can easily extract the task-related object mask and the operational workspace mask. Utilizing a hierarchical matching system based on SAM (Segment Anything Model), YOLOv8 (You only look once version 8), SIFT (Scale-Invariant Feature Transform), FLANN (Fast Library for Approximate Nearest Neighbors), and PCA (Principal Component Analysis), we can identify task-related objects and determine their configurations in new scenarios. Specifically, $\mathbf{M}=\bigcup_{k\in\mathbf{K}}M_{\text{SAM}}(\mathbf{H}\cdot\mathcal{G}(\mathbf{S}_{\text{therbligs}},k),\mathbf{B})$ represents object mask segmentation using the SAM model, while $\mathbf{Box}_{k}=M_{\text{YOLO}}(\mathbf{I}_{\text{new}},\text{ComputeArea}(\mathbf{m}_{k}))$ denotes bounding box detection through YOLOv8. The feature extraction is handled by $\mathbf{F}_{\text{new}}=\text{SIFT}(\mathbf{Box}_{k})$ using SIFT, feature matching is performed by $\mathbf{M}_{\text{match}}=\text{FLANN}(\mathbf{F}_{\text{new}},\mathbf{F}_{\text{ref}})$ using FLANN, and orientation extraction is achieved via $\mathbf{O}_{\text{new},k}=\text{PCA}(\mathbf{M}_{\text{match}})$ using PCA.
ActionREG excels in performing one-shot regression of object configurations, meaning it can accurately predict an object’s configuration in new and unseen environments using just a single example or image. This capability is vital for robotic systems operating in dynamic, real-world settings, ensuring robust and adaptable performance.
@@ III-D LAP-VC: LLM-Alignment Policy for Visual Correction
Expert demonstration may have some errors: robot end-effector grasping point may not always be perpendicular to the object itself. In addition, hand-eye calibration may also introduce some errors. These issues in real-world experiments may cause incorrect position estimation and affect our action registration results, especially in the grasp stage. To address these challenges, we introduced a novel method (Fig. 6) using LLM to correct these errors. By inputting our predicted points along with the scenario image into the LLM and utilizing a pre-built prompt, the LLM outputs corrected points. This method significantly reduces the impact of poor demonstrations and cumulative system errors, minimizing the need for highly accurate expert demonstrations.
Figure 6: Application of LAP-VC in robotic tasks. This process involves inputting predicted points along with the scenario image into the LLM. The pre-built prompt guides the LLM to output corrected points, compensating for errors due to factors such as imperfect grasping angles or hand-eye calibration inaccuracies.
@@ IV Experiments and Setup
To demonstrate the model’s generality, we developed the Omniverse Robot System (ORS) for convenient data collection. Offline data collection involved two individuals: one performing expert demonstrations and the other labeling the robot task’s status. The system collected data on robot joint angles, end effector pose, joint speed, force, and torque readings from the ’Robotiq’ sensor, task images, and real-time labels. We gathered data from six tasks (tool-pick&place, crossbeam-cutting, bricks-gluing, tissue-sweeping, surface-wiping, cup-pouring) with expert demonstrations. For online testing, a human expert performed new tasks in unseen scenarios, with ORS recording robot states and scenario images before and after the task. We tested five challenging tasks (board-rolling, foamblock-flipping, plate-scrubbing, spoon-tilting, and paper-stamping) to evaluate performance.
We utilized a UR5 robotic arm with an origami gripper, inspired by Liu et al. [24], for manipulation tasks. This setup improves adaptability and precision. Visual and force feedback were captured using a RealSense D435i camera and a RoboTiq force sensor. The setup ran on Ubuntu 22.04, with ROS2 Humble for component communication and MoveIt2 for motion planning. Development was conducted in VSCode and model training utilized an NVIDIA RTX 4090 GPU.
@@ V Results and Analysis
For offline training, we used robot states as input to segment therbligs. In this time-series segmentation task, our MGSF network outperforms other state-of-the-art methods. As shown in Table II, our network achieved the highest recall of 94.37%, surpassing traditional methods like Cascade SVM and deep learning models such as CNN-RNN and BiLSTM-Type3. In addition, we conducted an ablation study for our MGSF network to evaluate the impact of different components. The baseline recall of our backbone model (without fusion) was approximately 79. 77% with a transformer network. When we introduced the fusion mechanism, the recall improved to 84.29%. The addition of gate fusion mechanism boosted the recall to 90.92%. The final MGSF model, which incorporates gated fusion and meta-learning, achieved an recall of 94.37% (Table II).
@@
SM(ST)-SimScenario 13/50 (26%) 2/50 (4%) 11/50 (22%) 0/50 (0%) 15/50 (30%) 13.7%
BC(ST)-SimScenario 6/50 (12%) 0/50 (0%) 0/50 (0%) 2/50 (4%) 0/50 (0%) 2.7%
BC(MT)-SimScenario 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0%
Ours-SimScenario 48/50 (96%) 47/50 (94%) 46/50 (92%) 48/50 (96%) 47/50 (94%) 94.4%
SM(ST)-ComScenario 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0%
BC(ST)-ComScenario 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0%
BC(MT)-ComScenario 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0%
Ours-ComScenario 43/50 (86%) 42/50 (84%) 40/50 (80%) 34/50 (68%) 41/50 (82%) 80%
@@
Moreover, we also analyzed the recall of different therbligs in terms of various tasks (Fig. 7). Generally, surface wiping achieved the highest segmentation results (97.17%) with six diverse robot tasks, while tissue sweeping achieved the lowest results (93.88%). This discrepancy may be attributed to the complexity of tissue sweeping actions required to effectively put tissue into the dustpan. TLoad and Release achieved the lowest recall results, around 85.56% and 83.41% respectively. However, it is important to note that even though the release recall is relatively lower, this does not significantly affect the overall results. This is because, after intricate manipulation operations, the force-torque sensors may be affected by drift.
@@
Figure 7: Therblig segmentation recall for different robot tasks. It shows the recall rates for various therbligs across multiple robotic tasks, highlighting the performance of our segmentation network. For picka and place task, data points represented with an underscore indicate that the data is none with use.
@@
Furthermore, as shown in Fig. 8, the LAP-VC system consistently achieves high alignment performance scores, averaging 0.896 across various tasks, outperforming traditional methods such as KNN, SIFT, ORB, AKAZE, FAST, and BRISK. Its performance is only slightly lower than that of human experts manually performing alignment. Specifically, for the Stamp task, the LAP-VC system achieved a score of 0.84. This lower score is attributed to the relatively small size of the stamp. In contrast, the Sponge task achieved a higher score of 0.94 due to the sponge’s relatively simple and uniform structure.
For the results of task execution, Table III showed that our system can achieve promising performance in terms of simple scenario (SimScenario) and complex scenario (ComScenario) with different new tasks.
Figure 8: Performance comparison across various methods and tasks.
@@
For our SimScenario, we only consider two task-related and unseen objects appearing in our vision area. For ComScenario, we add 3-6 unrelated and unseen objects together with our task-related objects to mimic the real-world cluttered environments. For SimScenario mode, our system achieved around 94.4% average success rate for five tasks, while our success rate decreased to 80% if we switched to ComScenario mode. We find that spoon-tilting has a lower success rate because it requires a more dynamic and complex trajectory while robot solvers easily get trapped by singularity. We compared our system against Single-Task State Machine (ST-SM), Single-Task Behavior Clone (ST-BC), and Multi-Task Behavior Clone (MT-BC) baselines. State machines, built for single tasks, use pre-built policies and vision features to generate trajectories. In SimScenario, state machines achieve some success with well-built policies and excellent vision features. However, different tasks require different policies, significantly increasing system complexity, as noted by Luo et al. [16]. We collected 250 expert demonstrations for BC training and found that MT-BC training is challenging with limited data. ST-BC showed converging loss but resulted in unstable trajectories and task failures with minor errors. Applying BC to ComScenario led to divergent training and unpredictable trajectories.
ACT and diffusion policies were not used as baselines for a few reasons. Our system performs one-shot tasks and operates in zero-shot scenarios without requiring extensive pre-training for scene objects, thus needing less data. In contrast, ACT approaches typically require large datasets. Additionally, our system handles long-horizon tasks involving comprehensive action sequences, while diffusion policies focus on discrete sub-tasks. Therefore, a direct comparison may not provide meaningful insights. This distinction highlights our system’s data efficiency and generalization.
@@ VI Conclusion and Future Work
In this paper, we presented a novel TBBF to enhance the understanding and execution of robotic tasks. By decomposing complex tasks into fundamental therbligs, our innovative approach provides a structured and interpretable representation of robotic actions. The MGSF network was employed for accurate therblig segmentation during the offline training stage, while ActionREG efficiently facilitated precise action registration and integration of therbligs with object configurations during the online testing stage.
Our experimental results demonstrate the effectiveness of our framework, showcasing high recall in therblig segmentation and robust performance in real-world robot task execution. Specifically, we achieved results with 94.37% recall in therblig segmentation and impressive successful execution rates of 94.4% for new and long-horizon tasks in simple scenarios, and 80% in complex scenarios. However, some limitations should be addressed in future work. Firstly, our model relies heavily on standard expert demonstrations, which must follow specific protocols. It struggles with extremely noisy or unprofessional demonstrations. Secondly, our model, like many existing approaches, is affected by object shadows and light reflections. Additionally, we focused on the generation of 2D object configurations and did not account for 3D object configurations, which could be improved with depth information.
Future research will focus on refining the TBBF to enhance its robustness and adaptability in more complex and dynamic environments. We will also incorporate more user data to train a more robust action segmentation network. Furthermore, we plan to extend the applicability of our methods to a broader range of robotics platforms, facilitating the transfer of robot knowledge from the UR5 to other platforms such as Panda or Kinova robots.

### v3

@@ III-C ActionREG: SAM-driven action registration network
A cornerstone of our TBBF is the ActionREG, designed for reasoning context information and configuration (Fig. 2). Directly using SAM with geometric masks to predict object points can be unstable in cluttered environments without prior information. Instead, ActionREG integrates therbligs’ prior knowledge into the SAM model, enabling accurate reasoning about the objects involved in robotic tasks. This integration guides the model to better understand and predict object configurations within a task-specific context, ensuring reliable performance in complex scenarios.
Algorithm 2 ActionREG and Trajectory Generalization
@@
0:  New Trajectory based on new layouts $\mathbf{T}_{\text{new}}(\mathbf{I}_{\text{new}})$
@@
  $\mathbf{P}_{\text{demo}},\mathbf{O}_{\text{demo}}=\mathcal{P}_{\text{demo}}(\mathbf{S}_{\text{therbligs}},\mathbf{D})$
@@
  $\mathbf{\Delta}_{\text{transform}}=\mathcal{T}(\mathbf{P}_{\text{new}},\mathbf{P}_{\text{demo}})$
  $\mathbf{T}_{\text{new}}=\mathcal{A}(\mathcal{F}_{\text{demo}}(\mathbf{S}_{\text{therbligs}},\mathbf{D}),\mathbf{\Delta}_{\text{transform}})$
  return  $\mathbf{T}_{\text{new}}$
Through ActionREG, we efficiently extract task-related object masks and workspace configurations. The process starts with object mask segmentation via the Segment Anything Model (SAM), denoted as $M_{\text{SAM}}$. YOLOv8, represented as $M_{\text{YOLO}}$, then detects bounding boxes $\mathbf{Box}_{k}$ for each object. Features are extracted using SIFT, denoted by $\mathbf{F}_{\text{new}}$ for new images and $\mathbf{F}_{\text{ref}}$ for reference images. Feature matching is performed by FLANN, denoted as $\mathbf{M}_{\text{match}}$, and object orientations are determined using PCA, denoted by $\mathbf{O}_{\text{new},k}$. The transformation $\mathbf{\Delta}_{\text{transform}}$ is calculated using the function $\mathcal{T}$, which maps object positions from the demonstration ($\mathbf{P}_{\text{demo}}$) to the new environment ($\mathbf{P}_{\text{new}}$). The function $\mathcal{F}$ extracts the demonstration trajectory, and the new trajectory $\mathbf{T}_{\text{new}}$ is generated by applying the transformation $\mathbf{\Delta}_{\text{transform}}$ through the function $\mathcal{A}$. Then it can generalize to new configurations.
@@ III-D LAP-VC: LLM-Alignment Policy for Visual Correction
Expert demonstrations can have errors, such as the robot’s end-effector not grasping perpendicularly, and hand-eye calibration inaccuracies. These issues can lead to incorrect position estimations, especially during the grasp stage, impacting action registration. To mitigate these challenges, we developed a novel method (Fig. 6) that leverages LLM (GPT-4) for error correction. By feeding the predicted points and scenario image into the LLM with a pre-built prompt, the LLM provides corrected points. This approach minimizes the effects of imperfect demonstrations and system errors, reducing the reliance on highly accurate demonstrations.
Figure 6: Application of LAP-VC in robotic tasks. The pre-built prompt guides the LLM to process error points with scenario image and output corrected points, compensating for errors.
@@ IV Experiments and Setup
Offline data collection involved two individuals: one performing expert demonstrations and the other labeling the robot task’s status. We gathered data from six tasks: tool pick-and-place, crossbeam cutting, bricks gluing, tissue sweeping, surface wiping, and cup pouring. These tasks were selected for their well-defined and repeatable patterns, covering a range of basic actions such as grasping, cutting, sweeping, wiping, and pouring. These tasks provided the model with diverse experiences to effectively learn essential behaviors. The offline training system utilized 52 groups of demonstrations per task, amounting to a total of 312 demonstrations. For each demonstration, we collected trajectory data over a 60-second timeframe (10 Hz), resulting in 600 timesteps. Each sample contained 26-dimensional data, including joint angles, joint speeds, end-effector poses and orientation, forces, and torques. Thus, the dataset for each demonstration was structured as a 600 × 26 matrix. The offline training set included 52 demonstrations (600 timesteps, 26 features each) and was split into 60% training, 20% validation, and 20% testing. For online testing, a human expert performed new tasks in previously unseen scenarios. OSS recorded robot states and captured scenario images before and after each task. Five challenging tasks were used for testing: board rolling, foam block flipping, plate scrubbing, spoon tilting, and paper stamping. These tasks were selected to introduce novel movements, object interactions, and force dynamics which do not present in the training data.
@@ V Results and Analysis
For offline training, we used robot states to segment therbligs. Our MGSF network outperforms state-of-the-art methods in this time-series segmentation task. As shown in Table II, our network achieved an average recall of 94.37% across 20 random seeds, surpassing methods such as TCNs, Reformer and the LLM-based BERT model. We also manually designed threshold-based methods; however, they were extremely time-consuming to configure, lacked generalizability across tasks, and delivered poor performance overall. In addition, we conducted an ablation study for MGSF network to evaluate the impact of different components. The baseline recall of our backbone model (without fusion) was 79.77% (transformer). Introducing fusion mechanism increased the recall to 84.29%, and adding a gate fusion mechanism further boosted it to 90.92% (Table II). Our dataset ablation study tested 30 to 312 demonstrations across six tasks. The model stabilized around 210-270 demos (35-45 per task), with no significant improvements at 312 demos, highlighting MGSF’s data efficiency and rapid convergence.
@@
Ablation dataset 30 demos 90 demos 150 demos 210 demos 270 demos 312 demos (Ours)
MGSF (General Recall) $63.91\pm 7.89$ $76.06\pm 3.45$ $87.72\pm 2.26$ $92.46\pm 0.99$ $94.45\pm 0.27$ $94.37\pm 0.59$
MGSF (General precision) $60.10\pm 8.87$ $78.01\pm 2.72$ $87.88\pm 1.97$ $92.47\pm 1.01$ $94.50\pm 0.27$ $94.36\pm 0.60$
MGSF (F1-score) $59.15\pm 11.37$ $76.49\pm 3.14$ $87.70\pm 2.20$ $92.45\pm 1.00$ $94.46\pm 0.28$ $94.36\pm 0.60$
@@
SM + ST + SimScenario 13/50 (26%) 2/50 (4%) 11/50 (22%) 0/50 (0%) 15/50 (30%) 13.7%
BC + ST + SimScenario + one shot 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0%
BC + ST + SimScenario + 100 shots 6/50 (12%) 0/50 (0%) 0/50 (0%) 2/50 (4%) 0/50 (0%) 2.7%
BC + MT + SimScenario + one shot 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0%
BC + MT + SimScenario + 100 shots 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0%
LLM(TBBF) + MT + SimScenario + one shot 34/50 (68%) 28/50 (56%) 33/50 (66%) 31/50 (62%) 35/50 (70%) 64.4%
MGSF(TBBF) + MT + SimScenario + one shot (ours) 48/50 (96%) 47/50 (94%) 46/50 (92%) 48/50 (96%) 47/50 (94%) 94.4%
SM + ST + ComScenario 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0%
BC + ST + ComScenario + one shot 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0%
BC + ST + ComScenario + 100 shots 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0%
BC + MT + ComScenario + one shot 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0%
BC + MT + ComScenario + 100 shots 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0%
LLM(TBBF) + MT + ComScenario + one shot 29/50 (58%) 24/50 (48%) 30/50 (60%) 26/50 (52%) 32/50 (64%) 56.4%
MGSF(TBBF) + MT + ComScenario + one shot (ours) 43/50 (86%) 42/50 (84%) 40/50 (80%) 34/50 (68%) 41/50 (82%) 80%
Failure Case Analysis Therblig Segmentation Action Registration Context Matching Trajectory Planning Others Total
Ours + SimScenario 5/250 (2%) 2/250 (0.8%) 4/250 (1.6%) 2/250 (0.8%) 1/250 (0.4%) 5.6%
Ours + ComScenario 8/250 (3.2%) 11/250 (4.4%) 9/250 (3.6%) 13/250 (5.2%) 9/250 (3.6%) 20%
* SM means state machine methods, BC means behavior clone (shallow CNN based), LLM means large language model (BERT based), SimScenario means environment only contain task related objects (usually three objects), ComScenario mean environment contain many irrelevant objects (usually seven to ten objects), one shot means only provide one demonstration data (one image, one kinematic trajectory), 100 shots means providing input 100-demonstration data as input. Others in the failure study means factors such as demonstration failure, LAP-VC failure, potential collision, system crash.
@@
Moreover, we also analyzed the recall of different therbligs in terms of various tasks (Fig. 7). Note that these results are based on a single random seed and may slightly differ from the general recall results. Generally, surface wiping achieved the highest segmentation results (97.17%) with six diverse robot tasks, while tissue-sweeping achieved the lowest results (93.88%). This discrepancy may be attributed to the complexity of tissue-sweeping actions required to effectively put tissue into the dustpan. TLoad and Release achieved the lowest recall results, around 85.56% and 83.41% respectively. The lower release recall, caused by force-torque sensor drift, can occur after intricate manipulation operations.
@@
Figure 7: Therblig segmentation recall for different robot tasks. For pick and place task, data points represented with an underscore indicate that the data is none with use.
@@
Furthermore, the LAP-VC system (Fig. 8) consistently achieves high alignment performance scores, averaging 0.896 across various tasks, outperforming traditional methods such as KNN, SIFT, ORB, AKAZE, FAST, and BRISK. Its performance is slightly lower than that of human experts manually performing alignment. Specifically, for the Stamp task, the LAP-VC system achieved a score of 0.84. This lower score is attributed to the relatively small size of the stamp. In contrast, the Sponge task achieved a higher score of 0.94 due to sponge’s relatively simple and uniform structure.
Figure 8: Alignment performance comparison across various methods.
@@
For the results of task execution, Table III showed that our system can achieve promising performance for simple scenario (SimScenario) and complex scenario (ComScenario) with multi tasks and one shot data. For the SimScenario, we only consider two task-related and unseen objects appearing in image. For ComScenario, we add 3-6 unrelated and unseen objects together with our task-related objects to mimic the real-world cluttered environments. For SimScenario mode, our system achieved around 94.4% average success rate for five tasks, while our success rate decreased to 80% if we switched to ComScenario mode. We find that spoon-tilting has a lower success rate because it requires a more dynamic trajectory while robot solvers are easily trapped by singularity. We compared our system with State Machine, Behavior Clone and an LLM-based baseline. State machines, designed for single tasks, achieved some success in SimScenario with well-designed policies but struggled with increased complexity across tasks. For Behavior Cloning, both one-shot and 100-shot training were tested. Only single task and SimScenario with 100-shot works but most produced unstable trajectories prone to failure. The LLM baseline (with BERT) outperformed others when integrated into our one-shot TBBF but lagged behind our system due to lower therblig segmentation accuracy, demonstrating the critical impact of therblig segmentation module on task execution.
The failure case analysis underscores the interpretability of our TBBF system by identifying issues in specific modules. In SimScenario, the system demonstrates robustness with a low failure rate of 5.6%, mainly due to therblig segmentation (2%) and context matching (1.6%). Even in ComScenario, our system maintains good performance with a failure rate of 20%, where trajectory planning (5.2%) and action registration (4.4%) are the primary areas for improvement. This indicates that while our system performs reliably in both structured and complex environments, the ability to pinpoint module-specific issues allows us to enhance performance.
@@ VI Conclusion and Future Work
We presented a novel Therblig-Based Backbone Framework (TBBF) that enhances the understanding and execution of robotic tasks by decomposing complex tasks into fundamental therbligs, which serves as the core architecture enabling all key modules in our system. This therblig backbone allows for improved data efficiency, interpretability, and generalization by providing a structured representation upon which our modules operate. Our experimental results demonstrate the effectiveness of our framework, showcasing high recall in therblig segmentation and robust performance in real-world robot task execution. We achieved results with 94.37% recall in therblig segmentation and impressive successful execution rates of 94.4% for new and long-horizon tasks in simple scenarios, and 80% in complex scenarios.
However, some limitations should be addressed in future work. Firstly, our offline training dataset is relatively small. We plan to use more efficient data collection methods to build a larger and more diverse dataset. Additionally, we focused on 2D object configurations and plan to extend our approach to 3D configurations while incorporating geometric constraints to prevent collisions in complex environments. We also plan to conduct a more detailed analysis of the therblig backbone in future work, focusing on how each component impacts the overall success rate and comparing its performance with manually designed action primitives. Furthermore, we plan to deploy a local LLM model, such as LLaMa3, to reduce latency and improve processing efficiency.